rowansgranny:

You want to buy an item that costs $100. Which is the most cost-effective choice for buying the item?

Vocaloid:

answer choices a. using a paid membership card to buy it at a 10 percent discount. b. buying it online at a 10 percent discount with a $5 shipping charge. c. buying it at a 10 percent discount without sales tax.

Vocaloid:

think about the total cost of each option a) you get a 10% discount on the item, so 100 - 0.10(100) = 100 - 10 = 90 plus the membership fee (which is not stated) b) 10% discount on the item (bringing us down to 90) and then a 5 dollar shipping charge (bringing us up to $95) c) 10% discount without sales tax, so once again, we're at $90.